OpenStudy (precal):

At a company, there are "n" more males employees than female employees. If there are "k" male employees at a company, what fraction of the employees are male, in terms of n and k?

OpenStudy (akashdeepdeb):

You can use POE and also Plug-In here! Options there? :) Or do you want the algebraic method? :D

OpenStudy (precal):

yes I do have multiple choice I can list them

OpenStudy (akashdeepdeb):

I can give you the answer without the options! Do you want the actual algebraic method?

OpenStudy (precal):

sure

OpenStudy (akashdeepdeb):

there are "n" more males employees than female employees. If there are "k" male employees at a company, what fraction of the employees are male, in terms of n and k? Forget the 'In terms of n and k' --> It gets confusing otherwise. :D So the ACTUAL QUESTION is. there are "n" more males employees than female employees. If there are "k" male employees at a company, what fraction of the employees are male? Answer: male employees = k if males are n more, we can say that female emps. are n less. So female = k-n So Male emps. = male/total = k/(k + k-n) = k/(2k -n) Got it? :)

OpenStudy (akashdeepdeb):

Is there an answer like that? :/

OpenStudy (precal):

yes

OpenStudy (akashdeepdeb):

Then I suppose that is the answer, :D

OpenStudy (akashdeepdeb):

Can International students give the PSAT?

OpenStudy (precal):

Do you mean can international students take the PSAT?

OpenStudy (akashdeepdeb):

Yeah. :P

OpenStudy (precal):

I suppose if they are in high school in the US they can. Junior year is when students are competing for scholarships. Not sure if an international student can qualify. Here is the link if you would like to read up on it.

OpenStudy (precal):

www.nationalmerit.org

OpenStudy (akashdeepdeb):

I'll be applying to US universities next year! I want to go to the USA :D

OpenStudy (akashdeepdeb):

Do you live in the US?

OpenStudy (precal):

yes but PSAT are for high school students not college students

OpenStudy (precal):

You should do fine at the university level.

OpenStudy (akashdeepdeb):

I am in High school! I am 15 years old.

OpenStudy (precal):

You will still do fine.

OpenStudy (akashdeepdeb):

:D

OpenStudy (akashdeepdeb):

Where do you live in the US?

OpenStudy (precal):

in the south